40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8816     Entry time: Tue Jul 9 23:27:17 2013
Author: Koji 
Type: Summary 
Category: LSC 
Subject: MICH: ITMX/Y <=> PRM/BS 

The MICH actuation with PRM/BS was investigated again.

(ITMX -1 / ITMY +1) is equivalent to (PRM -0.267 and BS +0.50).


- PRMIsb was locked with REFL33I&AS55Q.

- Using the locking module in the LSC model, actuate ITMX (-1) and ITMY (+1) at 580.1Hz. Note that the notch filters in the MICH/PRCL servos were on.

- Look at the peak in the AS55Q spectrum. Tune the BS element in the output matrix of the lock-in to minimize the peak height.
=> The peak was minimized at BS = -0.50.

- Look at the peak in the REFL33I spectrum. Tune the PRM element in the output matrix of the lock-in to minimize the peak height.
=> The peak was minimized at PRM = +0.267

- These measurement leads to the conclusion mentioned above.

ELOG V3.1.3-